# Magnetic correlations and superconducting pairing near higher-order Van Hove singularities

### v1

@@ III Results and Discussion
We begin by setting the hopping parameters $t=1.00$, $t^{\prime}=0.10$, and Hubbard interaction $U=3.0|t|$, while varying TNN hopping $t^{\prime\prime}$ across different electron fillings $\langle n\rangle=0.50$, 0.75, 0.81, and 1.00. Our study focus on the influence of $t^{\prime\prime}$ on spin susceptibility. The results demonstrate distinct behaviors in spin susceptibility between two filling regions separated by the HOVH singularity. Notably, a crossover from FM to AFM correlations emerges at $\langle n\rangle=0.75$, as shown in Fig. 3.
Figure 3: The relationship between the spin susceptibility and $\mathbf{q}$ for different TNN hopping integrals $t^{\prime\prime}$ at fixed $U=3.0|t|$ and $\beta=6$. (a) $\langle n\rangle=0.50$; (b) $\langle n\rangle=0.75$; (c) $\langle n\rangle=0.81$; (d) $\langle n\rangle=1.00$.
Specifically, in the filling region below the HOVH singularity (i.e., $\langle n\rangle=0.50$), the system predominantly exhibits FM behavior, which is significantly enhanced with increasing $t^{\prime\prime}$. Near the HOVH filling at $\langle n\rangle=0.75$, FM and AFM correlations compete: $t^{\prime\prime}$ suppresses FM fluctuations at the $M$-point while enhancing it at the $\Gamma$-points. In contrast, within the filling range from HOVH to half-filling ($\langle n\rangle=1.0$), the system primarily displays AFM correlations that are suppressed by increasing $t^{\prime\prime}$. These results provide further support for previous findings.
@@
The results presented in Fig. 4 reveal that the $f_{n}$-wave pairing symmetry dominates over other symmetries, followed by the $d+id$-wave. As the electron filling $\langle n\rangle$ decreases from $0.973$ to $0.813$, the average $f_{n}$-wave pairing correlation is significantly enhanced. In contrast, when $\langle n\rangle$ deviates further from $0.813$, the $f_{n}$-wave correlations are suppressed. This underscores the pivotal role of the HOVH singularity in enhancing pairing correlations.
@@
Figure 5: Pairing correlation functions for different pairing symmetries with (a) $t^{\prime}=0$; (b) $t^{\prime}=0.10$; (c) $t^{\prime}=0.20$; (d) $t^{\prime}=0.30$.
@@
To investigate the influence of the NNN hopping $t^{\prime}$ on long-range superconducting pairing correlations, we fix the electron filling to $\langle n\rangle=0.813$, with hopping parameters $t=1.00$ and $t^{\prime\prime}=0.15$, while systematically varying $t^{\prime}$ from 0 to 0.30. As shown in Fig. 5, both the $f_{n}$-wave and $d+id$-wave pairing correlation functions exhibit a monotonic suppression as $t^{\prime}$ increases.
@@
A similar analysis of the effect of $t^{\prime\prime}$ on pairing correlations is performed with $\langle n\rangle=0.813$, $t=1.00$, and $t^{\prime}=0.20$, while $t^{\prime\prime}$ ranges from 0 to 0.30 in Fig. 6. Here, the $f_{n}$-wave remains dominant, and at $t^{\prime\prime}=0.15$, the superconducting pairing correlations exhibit an unexpected strengthening due to the HOVH effect.
@@
The observed transition from FM to AFM correlations across the HOVH singularity, coupled with the dominant $f_{n}$-wave superconducting pairing near $\langle n\rangle$=0.813, reveals a profound interplay between magnetic fluctuations and unconventional superconductivity. Below the HOVH filling $(\langle n\rangle<0.75$), the enhanced FM correlations with increasing $t^{\prime\prime}$ suggest a Stoner-type instability that typically competes with spin-singlet pairing. However, the emergence of robust $f_{n}$-wave superconductivity near $\langle n\rangle$=0.813, precisely where FM and AFM correlations compete, implies a complex mediation mechanism: while strong FM fluctuations would normally suppress conventional superconductivity, the proximity to the HOVH singularity may enable an unconventional pairing channel through the unique momentum-space structure of spin fluctuations near the HOVH point. These results collectively suggest that the HOVH singularity not only enhances pairing correlations through its singular density of states but also creates a unique magnetic environment where competing FM/AFM fluctuations may actively mediate unconventional superconductivity.
@@
Figure 6: Pairing correlation functions for different pairing symmetries with (a) $t^{\prime\prime}=0$; (b) $t^{\prime\prime}=0.05$; (c) $t^{\prime\prime}=0.15$; (d) $t^{\prime\prime}=0.30$.
@@
The influence of $t^{\prime\prime}$ on both FM and AFM correlations, alongside the non-monotonic enhancement of $f_{n}$-wave pairing at $t^{\prime\prime}$=0.15, further suggests that TNN hopping tunes the system through a sweet spot where the HOVH-enhanced DOS cooperates with residual magnetic fluctuations to stabilize $f_{n}$-wave pairing. This behavior bears resemblance to the spin-fluctuation-mediated pairing scenarios, where competition between different magnetic modes can promote unconventional pairing symmetries. The dominant of $f_{n}$-wave pairing may be particularly favored by the anisotropic spin fluctuations emerging near the HOVH filling, as evidenced by the $t^{\prime\prime}$-dependent modulation of susceptibility at $\Gamma$, $K$, and $M$ points.
@@ IV Conclusion
In this work, we have systematically investigated the magnetic correlations and superconducting pairing near HOVH singularities by employing DQMC and CPMC methods. We observed a crossover from FM to AFM correlations in the vicinity of the HOVH filling, driven by the interplay between TNN hopping $t^{\prime\prime}$ and electronic DOS divergence. Below the HOVH, $t^{\prime\prime}$ amplifies FM fluctuations, while suppressing AFM correlations at higher fillings approaching half-filling. In the filling regime $\langle n\rangle\sim 0.8$, $f_{n}$-wave pairing symmetry dominates, with HOVH-induced power-law divergence in the DOS significantly enhancing superconducting correlations. Interestingly, while larger $t^{\prime}$ and $t^{\prime\prime}$ values generally suppress $f_{n}$-wave pairing, a critical $t^{\prime\prime}=0.15$ triggers an anomalous enhancement of pairing correlations at a fix $t^{\prime}$ due to the HOVH effect. These results not only advance our understanding of the competitive mechanisms between magnetic fluctuations and unconventional superconductivity in HOVH systems but also provide a theoretical respect for tailoring correlated electronic phases in graphene-derived quantum materials through strain, doping, or interaction engineering.
@@
This work was supported by NSFC (12474218 and 12088101) and Beijing Natural Science Foundation (No. 1242022 and 1252022). The numerical simulations in this work were performed at the HSCC of Beijing Normal University.

### v2

@@ III Results and Discussion
We begin by setting the hopping parameters $t=1.00$, $t^{\prime}=0.10$, and Hubbard interaction $U=3.0|t|$, while varying TNN hopping $t^{\prime\prime}$ across different electron fillings $\langle n\rangle=0.50$, 0.75, 0.81, and 1.00. Our study focus on the influence of $t^{\prime\prime}$ on spin susceptibility. The results demonstrate distinct behaviors in spin susceptibility between two filling regions separated by the HOVH singularity. Notably, a crossover from FM to AFM correlations emerges at $\langle n\rangle=0.75$, as shown in Fig. 3. To best illustrate the evolution of magnetic fluctuations with $t^{\prime\prime}$ within each filling regime, we employed different vertical scales for each subplot.
Figure 3: The relationship between the spin susceptibility and $\mathbf{q}$ for different TNN hopping integrals $t^{\prime\prime}$ at fixed $U=3.0|t|$ and $\beta=6$ when $L=4$. (a) $\langle n\rangle=0.50$. (b) $\langle n\rangle=0.75$. (c) $\langle n\rangle=0.81$. (d) $\langle n\rangle=1.00$.
Specifically, in the filling region below the HOVH singularity (i.e., $\langle n\rangle=0.50$), the system predominantly exhibits FM behavior, which is significantly enhanced with increasing $t^{\prime\prime}$. Near the HOVH filling at $\langle n\rangle=0.75$, FM and AFM correlations compete: $t^{\prime\prime}$ suppresses fluctuations at the $M$ point while enhancing it at the $\Gamma$ points. In contrast, within the filling range from HOVH to half filling ($\langle n\rangle=1.0$), the system primarily displays AFM correlations that are suppressed by increasing $t^{\prime\prime}$. This momentum-space trend is further corroborated by real-space spin correlation analysis presented in the Appendix B. These results provide further support for previous findings.
@@
The results presented in Fig. 4 reveal that the $f_{n}$-wave pairing symmetry dominates over other symmetries, followed by the $d+id$ wave. As the electron filling $\langle n\rangle$ decreases from $0.973$ to $0.813$, the average $f_{n}$-wave pairing correlation is significantly enhanced. In contrast, when $\langle n\rangle$ deviates further from $0.813$, the $f_{n}$-wave correlations are suppressed. This underscores the pivotal role of the HOVH singularity in enhancing pairing correlations.
@@
Figure 5: Pairing correlation functions for different pairing symmetries with (a) $t^{\prime}=0$, (b) $t^{\prime}=0.10$, (c) $t^{\prime}=0.20$, and (d) $t^{\prime}=0.30$.
@@
To investigate the influence of the NNN hopping $t^{\prime}$ on long-range superconducting pairing correlations, we fix the electron filling to $\langle n\rangle=0.813$, with hopping parameters $t=1.00$ and $t^{\prime\prime}=0.15$, while systematically varying $t^{\prime}$ from 0 to 0.30. As shown in Fig. 5, both the $f_{n}$ wave and $d+id$ wave pairing correlation functions exhibit a monotonic suppression as $t^{\prime}$ increases.
@@
A similar analysis of the effect of $t^{\prime\prime}$ on pairing correlations is performed with $\langle n\rangle=0.813$, $t=1.00$, and $t^{\prime}=0.20$, while $t^{\prime\prime}$ ranges from 0 to 0.30 in Fig. 6. Here, the $f_{n}$ wave remains dominant, and at $t^{\prime\prime}=0.15$ the superconducting pairing correlations exhibit an unexpected strengthening due to the HOVH effect. Such a $t^{\prime\prime}$ value may be accessible in engineered systems through strain, doping, or substrate interactions[5, 27].
@@
The observed transition from FM to AFM correlations across the HOVH singularity, coupled with the dominant $f_{n}$-wave superconducting pairing near $\left\langle n\right\rangle$=0.813, reveals a profound interplay between magnetic fluctuations and unconventional superconductivity. Below the HOVH filling $(\left\langle n\right\rangle<0.75$), the enhanced FM correlations with increasing $t^{\prime\prime}$ suggest a Stoner-type instability that typically competes with spin-singlet pairing. However, the emergence of robust $f_{n}$-wave superconductivity near $\left\langle n\right\rangle$=0.813, precisely where FM and AFM correlations compete, implies a complex mediation mechanism: while strong FM fluctuations would normally suppress conventional superconductivity, the proximity to the HOVH singularity may enable an unconventional pairing channel through the unique momentum-space structure of spin fluctuations near the HOVH point. These results collectively suggest that the HOVH singularity not only enhances pairing correlations through its singular density of states but also creates a unique magnetic environment where competing FM/AFM fluctuations may actively mediate unconventional superconductivity.
@@
Figure 6: Pairing correlation functions for different pairing symmetries with (a) $t^{\prime\prime}=0$, (b) $t^{\prime\prime}=0.05$, (c) $t^{\prime\prime}=0.15$, and (d) $t^{\prime\prime}=0.30$.
@@
The influence of $t^{\prime\prime}$ on both FM and AFM correlations, alongside the nonmonotonic enhancement of $f_{n}$ wave pairing at $t^{\prime\prime}$=0.15, further suggests that TNN hopping tunes the system through a sweet spot where the HOVH-enhanced DOS cooperates with residual magnetic fluctuations to stabilize $f_{n}$-wave pairing. This behavior bears resemblance to the spin-fluctuation-mediated pairing scenarios, where competition between different magnetic modes can promote unconventional pairing symmetries. The dominance of $f_{n}$ wave pairing may be particularly favored by the anisotropic spin fluctuations emerging near the HOVH filling, as evidenced by the $t^{\prime\prime}$-dependent modulation of susceptibility at $\Gamma$, $K$, and $M$ points.
@@ IV Conclusion
In this work, we have systematically investigated the magnetic correlations and superconducting pairing near HOVH singularities by employing DQMC and CPMC methods. We observed a crossover from FM to AFM correlations in the vicinity of the HOVH filling, driven by the interplay between TNN hopping $t^{\prime\prime}$ and electronic DOS divergence. Below the HOVH, $t^{\prime\prime}$ amplifies FM fluctuations, while suppressing AFM correlations at higher fillings approaching half filling. In the filling regime $\langle n\rangle\approx 0.8$, $f_{n}$-wave pairing symmetry dominates, with HOVH-induced power-law divergence in the DOS significantly enhancing superconducting correlations. Interestingly, while larger $t^{\prime}$ and $t^{\prime\prime}$ values generally suppress $f_{n}$-wave pairing, a critical $t^{\prime\prime}=0.15$ triggers an anomalous enhancement of pairing correlations at a fixed $t^{\prime}$ due to the HOVH effect. These results not only advance our understanding of the competitive mechanisms between magnetic fluctuations and unconventional superconductivity in HOVH systems but also provide a theoretical respect for tailoring correlated electronic phases in graphene-derived quantum materials through strain, doping, or interaction engineering.
V Acknowledgements
@@
This work was supported by National Natural Science Foundation of China (Grants No. 12474218 and No. 12088101) and Beijing Natural Science Foundation (Grants No. 1242022 and 1252022). The numerical simulations in this work were performed at the HSCC (High-performance Scientific Computing Center) of Beijing Normal University.
VI Data availability
The data that support the findingsof this article are openly available.[48]
Appendix A Appendix A: FINITE-SIZE ANALYSIS
To further verify the robustness of our results against finite-size effects, we performed additional large-scale simulations for system size $L=6$.
Fig. 8 compares the spin susceptibility at $\langle n\rangle=0.81$ between $L=5$ and $6$. The qualitative behavior remains identical across system sizes. This demonstrates the convergence of magnetic correlations with increasing system size.
Figure 8: The relationship between the spin susceptibility and $\mathbf{q}$ for different $t^{\prime\prime}$ at fixed $\langle n\rangle=0.81$, $U=3.0|t|$, and $\beta=6$. (a) $L=5$. (b) $L=6$.
To assess the finite-size effects on the superconducting pairing correlations, we present in Fig. 9 the pairing correlation functions for different pairing symmetries across different system sizes $L=4$ and $6$. For each size, we select the closed-shell filling closest to the HOVH singularity: $\langle n\rangle=0.881$ for $L=4$, and $\langle n\rangle=0.843$ for $L=6$.
Figure 9: Pairing correlation functions for different pairing symmetries at $t=1$, $t^{\prime}=0.20$, $t^{\prime}=0.15$ when (a) $L=4$ and (b) $L=6$.
A direct comparison across Figs. 9(a) and 9(b) reveals a consistent physical picture: the $f_{n}$ wave pairing symmetry dominates over other symmetries near the HOVH filling, regardless of the system size.
Furthermore, the anomalous enhancement of $f_{n}$ wave pairing at the critical third-nearest-neighbor hopping $t^{\prime\prime}=0.15$, as discussed in the main text, is also robustly reproduced in the larger $L=6$ system. This key feature is explicitly demonstrated in Fig. 10, which plots the $f_{n}$ wave pairing correlation at $\langle n\rangle=0.843$ for $L=6$ under varying $t^{\prime\prime}$.
Figure 10: Pairing correlation functions for different pairing symmetries at $L=6$ when $t=1,t^{\prime}=0.20$, and (a) $t^{\prime\prime}=0$, (b) $t^{\prime\prime}=0.05$, (c) $t^{\prime\prime}=0.15$, and (d) $t^{\prime\prime}=0.30$.
The consistent dominance of $f_{n}$-wave pairing and the preservation of the critical $t^{\prime\prime}$ effect across different system sizes confirm that our key findings are intrinsic properties of the system rather than finite-size artifacts. These comprehensive finite-size analyses substantially strengthen the conclusions presented in the main text.
Appendix B Appendix B: REAL-SPACE SPIN CORRELATION ANALYSIS
To complement the momentum‑space analysis of spin fluctuations presented in Fig. 3(d), we examine the equal‑time nearest‑neighbor spin correlation $S^{Z}=\langle S_{i}^{z}S_{j}^{z}\rangle$ for $U=3.0|t|$ as $\langle n\rangle$ approaches 1.00.
As shown in Fig. 11, the AFM correlation (negative $S_{\langle i,j\rangle}$) is systematically suppressed as $t^{\prime\prime}$ increases, while changes in $t^{\prime}$ show negligible effect at <n>=1.00<n>=1.00. This real‑space result provides direct verification of the corresponding trend observed in momentum space in Fig. 3(d): the reduction in $\chi_{s}(M)$ with increasing $t^{\prime\prime}$.
Figure 11: The evolution of the equal‑time nearest‑neighbor spin correlation with filling for different $t^{\prime}$ and $t^{\prime\prime}$ with $U=3.0|t|$ and $L=4$.